What is the difference between two mean A.M and G.M of 2 and 8.

Grade:11

1 Answers

Arun
5 years ago
Dear Bipin

A.M. = (2+8)/2 = 5
While 
G.M. = sqrt ( 2*8) = sqrt(16) = 4
Hence difference = 5 ~4 = 1
Regards
Arun (askIITians forum expert)